# Neutron star mountains supported by crustal lattice pressure

### v1

@@ Abstract
pubyear: 2023pagerange: Neutron star mountains supported by crustal lattice pressureNeutron star mountains supported by crustal lattice pressure
@@ 1 Introduction
Spinning non-axisymmetric neutron stars (NSs) are a potential source of detectable continuous gravitational waves (CGWs); see Jones & Riles (2024) for a recent review. In order to support a non-axisymmetric shape, some deforming force beyond the axisymmetric centrifugal one must act, so that the star has what is colloquially referred to as a mountain.
The two types of deforming force to consider are magnetic and elastic ones; see Glampedakis & Gualtieri (2018) for a review. In the case of magnetic mountains, the Lorentz forces due to a non-axisymmetric magnetic field support the mass distortion. The existence of magnetic fields in NSs seems clear, with the (external) magnetic field having been invoked since the discovery of the first pulsars to explain their pulsation mechanism and their gradual spin-down.
In the case of elastic mountains, it is elastic strains in the solid crust that are responsible for the deformations. There have been several studies of the maximum mountains that a neutron star’s crust can support, limited by the finite shear modulus and breaking strain of the crustal material (Ushomirsky et al., 2000; Owen, 2005; Haskell et al., 2006; Johnson-McDaniel & Owen, 2013; Gittins et al., 2020; Gittins & Andersson, 2021; Morales & Horowitz, 2022). However, in the case of elastic mountains, the cause of the non-axisymmetry is less clear. The changing centrifugal forces in a spinning-up or spinning-down star are necessarily axisymmetric, so cannot directly contribute to CGW emission. It is possible that crustal fractures in response to spin-up or spin-down may result in non-axisymmetric distortions (Fattoyev et al., 2018), but the few detailed analyses of this concern themselves with maximum rather than likely mountain sizes (Giliberti & Cambiotti, 2022; Gangwar & Jones, 2024).
@@
In an attempt to supply the required temperature asymmetry, Osborne & Jones (2020) made use of the fact that accreting NSs in LMXBs have (albeit weak) magnetic fields. Such fields make the thermal conductivity tensor anisotropic, with heat being conducted more easily along the magnetic field lines than perpendicular to the field lines. This anisotropy in thermal conductivity naturally produces a temperature asymmetry. The crust-only model of Osborne & Jones (2020) was recently extended to the entire star by Hutchins & Jones (2023), who found somewhat larger temperature perturbations for a given field strength.
@@ 2.1 Mass quadrupole moment from capture layer shifts
$Q_{\text{fid}}\approx 1.3\times 10^{36}\,\text{g cm}^{2}\,R^{4}_{6}\,\bigg{(}\frac{T}{10^{8}\,\text{K}}\biggr{)}\,\bigg{(}\frac{\delta T/T}{1\%}\biggr{)}\,\bigg{(}\frac{E_{\text{cap}}}{30\,\text{MeV}}\biggr{)}^{3}\,,$ (2)
where $E_{\text{cap}}$ is the threshold energy, equivalent to the electron chemical potential $\mu_{e}$ at the transition between one nuclear species and the next. The above formula implies that, for a canonical $10$ km neutron star with a crustal temperature $\sim 10^{8}$ K and temperature asymmetry $\delta T/T\sim 1\%$, a single electron capture layer with threshold energy $E_{\text{cap}}\sim 30$ MeV produces a mass quadrupole which can be up to 3 orders of magnitude smaller than what Bildsten (1998) estimated would be required to establish torque balance, depending on the accretion rate and spin period of the system (see below).
In their calculation, UCB assumed (partially) the composition of the accreted crust to follow the series of non-equilibrium reactions listed in Tables 1 and 2 of Haensel & Zdunik (1990a) (the values of $\mu_{e}$ at each capture layer being listed in the sixth column of Table 2 of Haensel & Zdunik 1990b). The maximum value of the electron chemical potential (and therefore threshold energy) calculated by Haensel & Zdunik (1990a, b) (denoted collectively HZ90 hereafter) was $\mu_{e}\equiv E_{\text{max}}^{\text{HZ90}}=43.69$ MeV, suggesting that none of these capture layers are, individually, likely capable of producing a $Q_{22}$ large enough to attain torque balance as determined by equation (1).
UCB did, however, note that the total mass quadrupole of the NS $Q_{\text{tot}}$ is (approximately) a linear sum of the quadrupole moments generated in each capture layer individually. Combining the available data for $E_{\text{cap}}$ given in HZ90 and Eq. (2), one finds that the total mass quadrupole assuming a temperature asymmetry $\delta T/T\sim 1\%$ is
@@
where the summation includes all 19 capture layers listed in Table 2 of Haensel & Zdunik (1990b). Due to the strong scaling with the threshold energy (which increases monotonically through the crust), contributions from deep capture layers can be up to three orders of magnitude larger than those from shallow layers, and thus it is the deeper layers which dominate the total mass quadrupole.
@@
Even without considering elastic effects, the above result suggests that capture layer shifts are only capable of determining the spin equilibrium of LMXBs which are spinning very fast ($\sim 600$ Hz). Also, as shown in Osborne & Jones (2020); Hutchins & Jones (2023) (and UCB to a lesser extent), temperature asymmetries close to the (assumed) percent level are only likely to persist in systems which are strongly accreting. On this basis, it is therefore not clear that sufficiently large quadrupoles can be generated in the full set of LMXBs.
When extrapolating to higher values of the threshold energy however, UCB showed that a single additional ‘artificial’ capture layer (i.e. with $E_{\text{cap}}>44$ MeV) added ad-hoc near the bottom of the crust could generate mass quadrupoles in excess of $10^{37}$ g cm2, but only if the threshold energy exceeds $\sim 80$ MeV. UCB then considered the possibility of there being multiple high-threshold energy capture layers in the deep crust (see their Fig. 14). This led them to conclude that, collectively, these deeper layers can produce a mass quadrupole capable of matching Eq. (1), assuming temperature asymmetry at the level $\delta T/T\sim 1\%$.
@@
A more recent study by Gusakov & Chugunov (2020) (hereafter GC20) has even suggested that both the HZ90 and F+18 models are not in fact thermodynamically consistent, since they do not take into account the diffusion of unbound neutrons in the inner crust. When allowing for such diffusion, they found that the EoS of an accreted inner crust should actually be very close to that of cold-catalysed matter (see their Fig. 1). This is true not just for the pressure-density relation, but also for the composition, i.e. the run of atomic number $Z$ and mass number $A$. Composition tables for this so-called neutron Hydrostatic and Diffusion (nHD) model can be found in Potekhin et al. (2023) (Tables 1 and 2), where electron capture takes place only as far as up to neutron drip ($\sim 4\times 10^{11}$ g cm-3), and that $E_{\text{max}}^{\text{GC20}}\approx 25.5$ MeV.
@@ 2.2 Expected GW emission from capture layer shifts
The GW torque balance scenario (1) provides a rough upper limit to assess the detectability of CGWs from spinning neutron stars. It is an upper limit since the well known and well studied interaction between the star and the accretion disk, mediated by the star’s magnetic field, may play an important role in determining equilibrium spin frequencies (Ghosh et al., 1977; Ghosh & Lamb, 1979a, b; White & Zhang, 1997; Andersson et al., 2005), and may even be sufficient to explain the lack of observed LMXBs rotating faster than $\sim 700$ Hz (e.g. Patruno et al. 2012).
However, current observations certainly cannot definitively attribute the observed spin-rates of accreting neutron stars to that of the magnetic spin-equilibrium model alone. Indeed, it was noted by Bhattacharyya & Chakrabarty (2017) that an external magnetic field $B\sim 10^{8}$ G (typical of LMXBs) would be insufficient to explain the spin-limit of systems which accrete transiently. Furthermore, a statistical analysis of the spin-distribution of LMXBs by Patruno et al. (2017) suggests evidence for two distinct sub-populations: the majority of NSs having an average spin frequency $\bar{\nu}\approx 300$ Hz, and a smaller number of NSs having an average spin frequency $\bar{\nu}\approx 575$ Hz. A curious aspect of this observation is that if there is indeed an additional braking mechanism in the system, then it must set in sharply once a star reaches a given spin rate. It just so happens that, in the case of a deformed (rotating) NS, the rate of angular momentum loss via GWs scales as a steep (fifth) power of the star’s spin frequency (Bildsten, 1998). Another statistical analysis by Gittins & Andersson (2019) also found that a qualitatively similar behaviour to the observed distribution of transiently accreting NSs can indeed be obtained from GW spin-equilibrium models. Though, it is worth pointing out that the different GW mechanisms that the authors considered – a permanent quadrupole, unstable r-modes, and thermal mountains (from capture layer shifts) – all produced almost indistinguishable distributions.
@@
For targeted searches for CGWs from known pulsars (where all parameters of the source are known to a sufficiently high accuracy), the factor $A$ is often taken to be $11.4$, corresponding to a specific S/N threshold that permits a single trial false alarm rate of 1% and false dismissal rate of 10% (Abbott et al., 2004). For other types of searches where less is assumed about the signal (e.g. blind all-sky searches), a higher threshold would need to be applied. We will use $A=11.4$ in the analysis here, i.e. we assume the phase evolution of the signal is known from electromagnetic observations.
@@
In the left-hand panel of Figure 1 we plot the GW strain as per equation (5) for the largest thermal mountains that could be formed in a number of known LMXBs (plotted as triangles) by the deepest capture layer (corresponding to $E_{\rm{cap}}^{\rm{max}}$) assuming different accreted equations of state (UCB: Ushomirsky et al. 2000, HZ90: Haensel & Zdunik 1990b, F+18: Fantina et al. 2018, GC20: Gusakov & Chugunov 2020) with fixed maximum temperature asymmetry $\delta T/T=1\%$. The quadrupole moment is computed using equation (2). The LMXBs we choose to model are given in Table 1, and are a selection of those considered in Haskell et al. (2015). Specifically, we consider systems which have had an active outburst duration $\Delta t$ longer than two years, and for which the spin period $\nu$, time-averaged accretion rates $\langle\dot{M}\rangle$, and distances $d$ are either known or have been estimated. Only sources which lie above a given noisecurve are considered detectable.
@@
It is clear from Figure 1 that, even with a 2 year observation time, systems with mountains formed from shifting capture layers in the outer crust ($E_{\rm{cap}}<30$ MeV) not only fail to approach the torque balance limit imposed by equation (1) (marked for each system by stars in the Figure), but also fall below the sensitivity curves of both current and next-generation instruments. Even within such an optimistic analysis (i.e. ignoring elastic effects and assuming $\delta T/T=1\%$), capture layers in the inner crust (in conflict with the inner crust composition computed by Gusakov & Chugunov 2020) must be present if such mountains are to be detected by either Einstein Telescope or Cosmic Explorer.
While capture layers in the deep crust with $E_{\rm{cap}}\geq 90$ MeV exceed the torque balance limit of each of the systems we consider, we reiterate again that the existence of such layers are not supported by any detailed calculations of the accreted equation of state. And although capture layers with $E_{\rm{cap}}\approx 69$ MeV (predicted in the realistic F+18 model) could, in principle, allow for detection of all of the systems considered here in next-generation detectors, we also note that temperature asymmetry at the percent level is unlikely to be realised in these systems on account of their low time-averaged accretion rates (Osborne & Jones, 2020; Hutchins & Jones, 2023).
The chosen fixed temperature asymmetry $\delta T/T=1\%$ is in reality only an upper limit, imposed by the fact that the temperature perturbations which displace the capture layers should result in lateral variations in the flux emanating from the top of the crust. It is thought that if $\delta T/T\gg 1\%$, the resultant ‘hot and cold spots’ would lead to detectable modulations in quiescent emission (Ushomirsky et al., 2000), which so far have yet to be observed.
@@
Source $\nu$ $d$ $\langle\dot{M}\rangle$ $\Delta t$ Ref.
@@
Table 1: Estimated spin frequency $\nu$, distance $d$, and time-averaged accretion rate $\langle\dot{M}\rangle$ of a number of low-mass X-ray binaries with observed outburst duration’s $\Delta t$ longer than 2 years. Sources marked (†) are nuclear powered pulsars, while HETE J1900.1–2455 is an accreting millisecond pulsar. Note that this is an adapted reproduction of Table 1 from Haskell et al. (2015).
Figure 1: Plots of signal strength $h_{0}$ (triangles/stars) and minimum detectable amplitudes $h_{0,\,\rm{det}}$ (curves) for various signal models and GW detectors. The minimum detectable amplitudes were obtained from equation (7) with $A=11.4$ and $T_{\rm obs}=2$ years. In the left panel we plot the torque balance limit as per equation (1) for the low-mass X-ray binaries listed in Table 1, and estimate the largest thermal mountain that can be created with an assumed maximum temperature asymmetry $\delta T/T=1\%$ for different accreted equations of state (UCB: Ushomirsky et al. 2000, HZ90: Haensel & Zdunik 1990b, F+18: Fantina et al. 2018, GC20: Gusakov & Chugunov 2020) using the fiducial estimate (2). In the right panel we show the effects on mountain sizes which arise from elastic readjustments of the crust to shifting deep (90 MeV) and shallow (25 MeV) capture layers. The bars indicate the range of uncertainty in $Q_{\rm{fid}}$ (and thus $h_{0}$) that arise due to ‘sinking penalties’.

### v2

# Crustal lattice pressure as a source of neutron star mountains

@@ Abstract
pubyear: 2023pagerange: Crustal lattice pressure as a source of neutron star mountainsCrustal lattice pressure as a source of neutron star mountains
@@ 1 Introduction
Spinning non-axisymmetric neutron stars (NSs) are a potential source of detectable continuous gravitational waves (CGWs); see Jones and Riles (2024) for a recent review. In order to support a non-axisymmetric shape, some deforming force beyond the axisymmetric centrifugal one must act, so that the star has what is colloquially referred to as a mountain.
The two types of deforming force to consider are magnetic and elastic ones; see Glampedakis and Gualtieri (2018) for a review. In the case of magnetic mountains, the Lorentz forces due to a non-axisymmetric magnetic field support the mass distortion. The existence of magnetic fields in NSs seems clear, with the (external) magnetic field having been invoked since the discovery of the first pulsars to explain their pulsation mechanism and their gradual spin-down.
In the case of elastic mountains, it is elastic strains in the solid crust that are responsible for the deformations. There have been several studies of the maximum mountains that a neutron star’s crust can support, limited by the finite shear modulus and breaking strain of the crustal material (Ushomirsky et al., 2000; Owen, 2005; Haskell et al., 2006; Johnson-McDaniel and Owen, 2013; Gittins et al., 2020; Gittins and Andersson, 2021; Morales and Horowitz, 2022). However, in the case of elastic mountains, the cause of the non-axisymmetry is less clear. The changing centrifugal forces in a spinning-up or spinning-down star are necessarily axisymmetric, so cannot directly contribute to CGW emission. It is possible that crustal fractures in response to spin-up or spin-down may result in non-axisymmetric distortions (Fattoyev et al., 2018), but the few detailed analyses of mountain formation following crustal fracture concern themselves with maximum rather than likely mountain sizes (Giliberti and Cambiotti, 2022; Gangwar and Jones, 2024).
@@
In an attempt to supply the required temperature asymmetry, Osborne and Jones (2020) made use of the fact that accreting NSs in LMXBs have (albeit weak) magnetic fields. Such fields make the thermal conductivity tensor anisotropic, with heat being conducted more easily along the magnetic field lines than perpendicular to the field lines. This anisotropy in thermal conductivity naturally produces a temperature asymmetry. The crust-only model of Osborne and Jones (2020) was recently extended to the entire star by Hutchins and Jones (2023), who found somewhat larger temperature perturbations for a given field strength.
@@ 2.1 Mass quadrupole moment from capture layer shifts
$Q_{\text{fid}}\approx 1.3\times 10^{36}\,\text{g cm}^{2}\,R^{4}_{6}\,\bigg(\frac{T}{10^{8}\,\text{K}}\biggr)\,\bigg(\frac{\delta T_{22}/T}{1\%}\biggr)\,\bigg(\frac{E_{\text{cap}}}{30\,\text{MeV}}\biggr)^{3}\,$, (2)
where $E_{\text{cap}}$ is the threshold energy, equivalent to the electron chemical potential $\mu_{e}$ at the transition between one nuclear species and the next. The above formula implies that, for a canonical $10$ km neutron star with a crustal temperature $\sim 10^{8}$ K and quadrupolar ($l=m=2$) temperature asymmetry $\delta T_{22}/T\sim 1\%$, a single electron capture layer with threshold energy $E_{\text{cap}}\sim 30$ MeV produces a mass quadrupole which can be up to 3 orders of magnitude smaller than what Bildsten (1998) estimated would be required to establish torque balance, depending on the accretion rate and spin period of the system (see below).
In their calculation, UCB assumed (partially) the composition of the accreted crust to follow the series of non-equilibrium reactions listed in Tables 1 and 2 of Haensel and Zdunik (1990b) (the values of $\mu_{e}$ at each capture layer being listed in the sixth column of Table 2 of Haensel and Zdunik 1990a). The maximum value of the electron chemical potential (and therefore threshold energy) calculated by Haensel and Zdunik (1990b, a) (denoted collectively HZ90 hereafter) was $\mu_{e}\equiv E_{\text{max}}^{\text{HZ90}}=43.69$ MeV, suggesting that none of these capture layers are, individually, likely capable of producing a $Q_{22}$ large enough to attain torque balance as determined by equation (1).
UCB did, however, note that the total mass quadrupole of the NS $Q_{\text{tot}}$ is (approximately) a linear sum of the quadrupole moments generated in each capture layer individually. Combining the available data for $E_{\text{cap}}$ given in HZ90 and Eq. (2), one finds that the total mass quadrupole assuming a temperature asymmetry $\delta T_{22}/T\sim 1\%$ is
@@
where the summation includes all 19 capture layers listed in Table 2 of Haensel and Zdunik (1990a). Due to the strong scaling with the threshold energy (which increases monotonically through the crust), contributions from deep capture layers can be up to three orders of magnitude larger than those from shallow layers, and thus it is the deeper layers which dominate the total mass quadrupole.
@@
Even without considering elastic effects, the above result suggests that capture layer shifts are only capable of determining the spin equilibrium of LMXBs which are spinning very fast ($\sim 600$ Hz). Also, as shown in Osborne and Jones (2020); Hutchins and Jones (2023) (and UCB to a lesser extent), temperature asymmetries close to the (assumed) percent level are only likely to persist in systems which are strongly accreting. On this basis, it is therefore not clear that sufficiently large quadrupoles can be generated in the full set of LMXBs.
When extrapolating to higher values of the threshold energy however, UCB showed that a single additional ‘artificial’ capture layer (i.e. with $E_{\text{cap}}>44$ MeV) added ad-hoc near the bottom of the crust could generate mass quadrupoles in excess of $10^{37}$ g cm2, but only if the threshold energy exceeds $\sim 80$ MeV. UCB then considered the possibility of there being multiple high-threshold energy capture layers in the deep crust (see their Fig. 14). This led them to conclude that, collectively, these deeper layers can produce a mass quadrupole capable of matching Eq. (1), assuming temperature asymmetry at the level $\delta T_{22}/T\sim 1\%$.
@@
A more recent study by Gusakov and Chugunov (2020) (hereafter GC20) has even suggested that both the HZ90 and F+18 models are not in fact thermodynamically consistent, since they do not take into account the diffusion of unbound neutrons in the inner crust. When allowing for such diffusion, they found that the EoS of an accreted inner crust should actually be very close to that of cold-catalysed matter (see their Fig. 1). This is true not just for the pressure-density relation, but also for the composition, i.e. the run of atomic number $Z$ and mass number $A$. Composition tables for this so-called neutron Hydrostatic and Diffusion (nHD) model can be found in Potekhin et al. (2023) (Tables 1 and 2), where electron capture takes place only as far as up to neutron drip ($\sim 4\times 10^{11}$ g cm-3), and that $E_{\text{max}}^{\text{GC20}}\approx 25.5$ MeV.
@@ 2.2 Expected GW emission from capture layer shifts
The GW torque balance scenario (1) provides a rough upper limit to assess the detectability of CGWs from spinning neutron stars. It is an upper limit since the well known and well studied interaction between the star and the accretion disk, mediated by the star’s magnetic field, may play an important role in determining equilibrium spin frequencies (Ghosh et al., 1977; Ghosh and Lamb, 1979a, b; White and Zhang, 1997; Andersson et al., 2005), and may even be sufficient to explain the lack of observed LMXBs rotating faster than $\sim 700$ Hz (e.g. Patruno et al. 2012).
However, current observations certainly cannot definitively attribute the observed spin-rates of accreting neutron stars to that of the magnetic spin-equilibrium model alone. Indeed, it was noted by Bhattacharyya and Chakrabarty (2017) that an external magnetic field $B\sim 10^{8}$ G (typical of LMXBs) would be insufficient to explain the spin-limit of systems which accrete transiently. Furthermore, a statistical analysis of the spin-distribution of LMXBs by Patruno et al. (2017) suggests evidence for two distinct sub-populations: the majority of NSs having an average spin frequency $\bar{\nu}\approx 300$ Hz, and a smaller number of NSs having an average spin frequency $\bar{\nu}\approx 575$ Hz. A curious aspect of this observation is that if there is indeed an additional braking mechanism in the system, then it must set in sharply once a star reaches a given spin rate. It just so happens that, in the case of a deformed (rotating) NS, the rate of angular momentum loss via GWs scales as a steep (fifth) power of the star’s spin frequency (Bildsten, 1998). Another statistical analysis by Gittins and Andersson (2019) also found that a qualitatively similar behaviour to the observed distribution of transiently accreting NSs can indeed be obtained from GW spin-equilibrium models. Though, it is worth pointing out that the different GW mechanisms that the authors considered – a permanent quadrupole, unstable r-modes, and thermal mountains (from capture layer shifts) – all produced almost indistinguishable distributions.
@@
For targeted searches for CGWs from known pulsars (where all parameters of the source are known to a sufficiently high accuracy), the factor $A$ is often taken to be $11.4$, corresponding to a specific S/N threshold that permits a single trial false alarm rate of 1% and false dismissal rate of 10% (Abbott and others, 2004). For other types of searches where less is assumed about the signal (e.g. blind all-sky searches), a higher threshold would need to be applied. We will use $A=11.4$ in the analysis here, i.e. we assume the phase evolution of the signal is known from electromagnetic observations.
@@
In the left-hand panel of Figure 1 we plot the GW strain as per equation (5) for the largest thermal mountains that could be formed in a number of known LMXBs (plotted as triangles) by the deepest capture layer (corresponding to $E_{\rm{cap}}^{\rm{max}}$) assuming different accreted equations of state (UCB: Ushomirsky et al. 2000, HZ90: Haensel and Zdunik 1990a, F+18: Fantina et al. 2018, GC20: Gusakov and Chugunov 2020) with fixed maximum temperature asymmetry $\delta T_{22}/T=1\%$. The quadrupole moment is computed using equation (2). The LMXBs we choose to model are given in Table 1, and are a selection of those considered in Haskell et al. (2015). Specifically, we consider systems which have had an active outburst duration $\delta T_{22}$ longer than two years, and for which the spin period $\nu$, time-averaged accretion rates $\langle\dot{M}\rangle$, and distances $d$ are either known or have been estimated. Only sources which lie above a given noisecurve are considered detectable.
@@
It is clear from Figure 1 that, even with a 2 year observation time, systems with mountains formed from shifting capture layers in the outer crust ($E_{\rm{cap}}<30$ MeV) not only fail to approach the torque balance limit imposed by equation (1) (marked for each system by stars in the Figure), but also fall below the sensitivity curves of both current and next-generation instruments. Even within such an optimistic analysis (i.e. ignoring elastic effects and assuming $\delta T_{22}/T=1\%$), capture layers in the inner crust (in conflict with the inner crust composition computed by Gusakov and Chugunov 2020) must be present if such mountains are to be detected by either Einstein Telescope or Cosmic Explorer.
While capture layers in the deep crust with $E_{\rm{cap}}\geq 90$ MeV exceed the torque balance limit of each of the systems we consider, we reiterate again that the existence of such layers are not supported by any detailed calculations of the accreted equation of state. And although capture layers with $E_{\rm{cap}}\approx 69$ MeV (predicted in the realistic F+18 model) could, in principle, allow for detection of all of the systems considered here in next-generation detectors, we also note that temperature asymmetry at the percent level is unlikely to be realised in these systems on account of their low time-averaged accretion rates (Osborne and Jones, 2020; Hutchins and Jones, 2023).
The chosen fixed temperature asymmetry $\delta T_{22}/T=1\%$ is in reality only an upper limit, imposed by the fact that the temperature perturbations which displace the capture layers should result in lateral variations in the flux emanating from the top of the crust. It is thought that if $\delta T_{22}/T\gg 1\%$, the resultant ‘hot and cold spots’ would lead to detectable modulations in quiescent emission (Ushomirsky et al., 2000), which so far have yet to be observed.
@@
Source $\nu$ $d$ $\langle\dot{M}\rangle$ $\delta T_{22}$ Ref.
@@
Table 1: Estimated spin frequency $\nu$, distance $d$, and time-averaged accretion rate $\langle\dot{M}\rangle$ of a number of low-mass X-ray binaries with observed outburst duration’s $\delta T_{22}$ longer than 2 years. Sources marked (†) are nuclear powered pulsars, while HETE J1900.1–2455 is an accreting millisecond pulsar. Note that this is an adapted reproduction of Table 1 from Haskell et al. (2015).
Figure 1: Plots of signal strength $h_{0}$ (triangles/stars) and minimum detectable amplitudes $h_{0,\,\rm{det}}$ (curves) for various signal models and GW detectors. The minimum detectable amplitudes were obtained from equation (7) with $A=11.4$ and $T_{\rm obs}=2$ years. In the left panel we plot the torque balance limit as per equation (1) for the low-mass X-ray binaries listed in Table 1, and estimate the largest thermal mountain that can be created with an assumed maximum temperature asymmetry $\delta T_{22}/T=1\%$ for different accreted equations of state (UCB: Ushomirsky et al. 2000, HZ90: Haensel and Zdunik 1990a, F+18: Fantina et al. 2018, GC20: Gusakov and Chugunov 2020) using the fiducial estimate (2). In the right panel we show the effects on mountain sizes which arise from elastic readjustments of the crust to shifting deep (90 MeV) and shallow (25 MeV) capture layers. The bars indicate the range of uncertainty in $Q_{\rm{fid}}$ (and thus $h_{0}$) that arise due to ‘sinking penalties’.